OpenStudy (anonymous):

how do i solve the equation 29 1/2 - 2 3/8

OpenStudy (anonymous):

convert 1/2 to eighths, 1/2 = 4/8 then do simple subtraction 29 4/8 - 2 3/8 = 27 1/8